Franklin and Tan collected an equal number of empty cans for an individual art assignment. Seeing that Tan needed some more cans for his design, Franklin gave him 96 cans. Tan had 5 times as many cans as Franklin after receiving the cans from Franklin. How many cans had Tan in the end?
1
Expert's answer
2020-04-07T09:42:13-0400

x is the initial number of cans each had

x + 96 = 5*(x - 96)

4x = 576

x = 144

y is the number of cans of Tan in the end

y = x + 96 = 240